# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ D.1 Linear GCN
In this analysis, we consider a simple but effective variant of GNNs, linear GCN, which is adapted from the classical SGC model wu2019simplifying .
@@
In the training process of Linear GCNs, we consider a general least squares problem shin2023efficient , which applies to both regression and binary classification tasks.
@@
D.2 Edge Influence on Model Weight
@@
For an arbitrary edge subset $\Delta\mathcal{E}$, the corresponding perturbation matrix $\boldsymbol{\Xi}^{\Delta\mathcal{E}}$ is defined as:
$\boldsymbol{\Xi}^{\Delta\mathcal{E}}:=\sum_{(v_{i},v_{j})\in\Delta\mathcal{E}}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}.$
For notation simplicity, we sometimes ignore $\Delta\mathcal{E}$ and use $\boldsymbol{\Xi}$ as a shorthand notation for $\boldsymbol{\Xi}$ in this paper.
@@
Let $\boldsymbol{\Xi}^{\Delta\mathcal{E}}\in\mathbb{R}^{n\times n}$ be a perturbation matrix as defined in Definition D.6. Let $\epsilon\in\mathbb{R}$ be a perturbation magnitude and $\mathbf{w}^{\star}(\epsilon)$ be the optimal model weight after perturbation. Considering an edge perturbation $\mathbf{C}(\epsilon):=\mathbf{C}+\epsilon\boldsymbol{\Xi}^{\Delta\mathcal{E}}$ on the original propagation matrix $\mathbf{C}$, the sensitivity of the model weight $\mathbf{w}^{\star}$ can be characterized by:
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star}):=\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Bigr{|}_{\epsilon=0}=-\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi}^{\Delta\mathcal{E}},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr{|}_{\epsilon=0},$
@@
For notation simplicity, we use $\boldsymbol{\Xi}$ as a shorthand notation for $\boldsymbol{\Xi}^{\Delta\mathcal{E}}$ in this proof. Since $\mathbf{w}^{\star}(\epsilon)$ is the minimizer of the perturbed loss function $\mathcal{L}(\mathbf{C}(\epsilon),\mathbf{X},\mathbf{w},\mathbf{y})$, we have:
$\mathbf{w}^{\star}(\epsilon)=\mathop{\arg\min}_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w},\mathbf{y}).$
@@
$\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star}(\epsilon),\mathbf{y})=0.$ (7)
@@
00 0 $=~{}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star}(\epsilon),\mathbf{y})$ (8)
$=~{}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star}+\Delta\mathbf{w},\mathbf{y})$
$=~{}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})+\nabla^{2}_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Delta\mathbf{w}$
$\quad\quad\quad+\epsilon\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr{|}_{\epsilon=0}+o(\lVert\Delta\mathbf{w}\rVert+\lvert\epsilon\rvert)$
@@
$\quad\quad\quad+\epsilon\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr{|}_{\epsilon=0}+o(\lVert\Delta\mathbf{w}\rVert+\lvert\epsilon\rvert)$
$=~{}\nabla^{2}_{\mathbf{w}}\mathcal{L}(\mathbf{C},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Delta\mathbf{w}+\epsilon\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr{|}_{\epsilon=0}$
$\quad\quad\quad+o(\lVert\Delta\mathbf{w}\rVert+\lvert\epsilon\rvert)$
@@
$\Delta\mathbf{w}=-\epsilon\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr{|}_{\epsilon=0}.$
@@
$\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Bigr{|}_{\epsilon=0}=-\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr{|}_{\epsilon=0}.$
@@
Given an arbitrary edge perturbation matrix $\boldsymbol{\Xi}$ as defined in Definition D.6, the influence function for the optimal model weight $\mathbf{w}^{\star}$ is:
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star}):=\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Bigr{|}_{\epsilon=0}=\mathbf{H}^{-1}(\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\mathbf{C}^{\top}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}).$
@@
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star}):=\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Biggr{|}_{\epsilon=0}=-\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Biggr{|}_{\epsilon=0},$
@@
Recalling the loss function in Definition D.4, when we introduce the perturbation $\mathbf{C}(\epsilon)=\mathbf{C}+\epsilon\boldsymbol{\Xi}$, the loss becomes
$\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w},\mathbf{y})=\frac{1}{2}\|\mathbf{y}-(\mathbf{C}+\epsilon\boldsymbol{\Xi})\mathbf{X}\mathbf{w}\|_{2}^{2}.$
@@
$\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w},\mathbf{y})=-\mathbf{X}^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})^{\to}\mathbf{P}\bigl{[}\mathbf{y}-(\mathbf{C}+\epsilon\boldsymbol{\Xi})\mathbf{X}\mathbf{w}\bigr{]}.$
@@
$\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})=-\mathbf{X}^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})^{\to}\mathbf{P}\bigl{[}\mathbf{y}-(\mathbf{C}+\epsilon\boldsymbol{\Xi})\mathbf{X}\mathbf{w}^{\star}\bigr{]}.$
@@
$\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})=$ $~{}-\underbrace{\mathbf{X}^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})^{\top}\mathbf{y}}_{:=\mathbf{T}_{1}}+\underbrace{\mathbf{X}^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})\mathbf{X}\mathbf{w}^{\star}}_{:=\mathbf{T}_{2}}$
@@
$\frac{\mathrm{d}\mathbf{T}_{1}}{\mathrm{d}\epsilon}=$ $~{}\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Bigl{[}-\mathbf{X}^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})^{\top}y\Bigr{]}$
$=$ $~{}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}y.$
@@
$\frac{\mathrm{d}\mathbf{T}_{2}}{\mathrm{d}\epsilon}=$ $~{}\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Bigl{[}\mathbf{X}^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})^{\top}(\mathbf{C}+\epsilon\boldsymbol{\Xi})\mathbf{X}\mathbf{w}^{\star}\Bigr{]}$
$=$ $~{}\mathbf{X}^{\top}(\mathbf{C}^{\top}\boldsymbol{\Xi}+\boldsymbol{\Xi}^{\top}\mathbf{C}+2\epsilon\boldsymbol{\Xi}^{\top}\boldsymbol{\Xi})\mathbf{X}\mathbf{w}^{\star}$
@@
$\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Biggr{|}_{\epsilon=0}=-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}+\mathbf{X}^{\top}(\mathbf{C}^{\top}\boldsymbol{\Xi}+\boldsymbol{\Xi}^{\top}\mathbf{C})\mathbf{X}\mathbf{w}^{\star}.$
@@
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star})$ $=-\mathbf{H}^{-1}\Biggl{(}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}+\mathbf{X}^{\top}(\mathbf{C}^{\top}\boldsymbol{\Xi}+\boldsymbol{\Xi}^{\top}\mathbf{C})\mathbf{X}\mathbf{w}^{\star}\Biggr{)}$
$=\mathbf{H}^{-1}\Bigl{(}\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr{)}.$
@@
D.3 Edge Influence on Model Output

### v2

@@ D.1. Linear GCN
In this analysis, we consider a simple but effective variant of GNNs, linear GCN, which is adapted from the classical SGC model (wu2019simplifying).
@@
In the training process of Linear GCNs, we consider a general least squares problem (shin2023efficient), which applies to both regression and binary classification tasks.
@@
D.2. Edge Influence on Model Weight
@@
For an arbitrary edge subset $\Delta\mathcal{E}$, the corresponding perturbation matrix $\bm{\Xi}^{\Delta\mathcal{E}}$ is defined as:
$\bm{\Xi}^{\Delta\mathcal{E}}:=\sum_{(v_{i},v_{j})\in\Delta\mathcal{E}}\mathbf{e}_{i}\mathbf{e}_{j}^{\top}.$
For notation simplicity, we sometimes ignore $\Delta\mathcal{E}$ and use $\bm{\Xi}$ as a shorthand notation for $\bm{\Xi}$ in this paper.
@@
Let $\bm{\Xi}^{\Delta\mathcal{E}}\in\mathbb{R}^{n\times n}$ be a perturbation matrix as defined in Definition D.6. Let $\epsilon\in\mathbb{R}$ be a perturbation magnitude and $\mathbf{w}^{\star}(\epsilon)$ be the optimal model weight after perturbation. Considering an edge perturbation $\mathbf{C}(\epsilon):=\mathbf{C}+\epsilon\bm{\Xi}^{\Delta\mathcal{E}}$ on the original propagation matrix $\mathbf{C}$, the sensitivity of the model weight $\mathbf{w}^{\star}$ can be characterized by:
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star}):=$ $~\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Bigr|_{\epsilon=0}$
$=$ $~-\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi}^{\Delta\mathcal{E}},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr|_{\epsilon=0},$
@@
For notation simplicity, we use $\bm{\Xi}$ as a shorthand notation for $\bm{\Xi}^{\Delta\mathcal{E}}$ in this proof. Since $\mathbf{w}^{\star}(\epsilon)$ is the minimizer of the perturbed loss function $\mathcal{L}(\mathbf{C}(\epsilon),\mathbf{X},\mathbf{w},\mathbf{y})$, we have:
$\mathbf{w}^{\star}(\epsilon)=\mathop{\arg\min}_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w},\mathbf{y}).$
@@
(7) $\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star}(\epsilon),\mathbf{y})=0.$
@@
(8) $0$ $=~\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star}(\epsilon),\mathbf{y})$
$=~\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star}+\Delta\mathbf{w},\mathbf{y})$
$=~\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})+\nabla^{2}_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Delta\mathbf{w}$
[…]
@@
[…]
$=~\nabla^{2}_{\mathbf{w}}\mathcal{L}(\mathbf{C},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Delta\mathbf{w}+\epsilon\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr|_{\epsilon=0}$
[…]
@@
$\Delta\mathbf{w}=-\epsilon\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr|_{\epsilon=0}.$
@@
$\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Bigr|_{\epsilon=0}=-\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Bigr|_{\epsilon=0}.$
@@
Given an arbitrary edge perturbation matrix $\bm{\Xi}$ as defined in Definition D.6, the influence function for the optimal model weight $\mathbf{w}^{\star}$ is:
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star}):=$ $~\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Bigr|_{\epsilon=0}$
$=$ $~\mathbf{H}^{-1}(\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\mathbf{C}^{\top}\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}).$
@@
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star}):=\frac{\mathrm{d}\mathbf{w}^{\star}(\epsilon)}{\mathrm{d}\epsilon}\Biggr|_{\epsilon=0}=-\mathbf{H}^{-1}\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Biggr|_{\epsilon=0},$
@@
Recalling the loss function in Definition D.4, when we introduce the perturbation $\mathbf{C}(\epsilon)=\mathbf{C}+\epsilon\bm{\Xi}$, the loss becomes
$\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w},\mathbf{y})=\frac{1}{2}\|\mathbf{y}-(\mathbf{C}+\epsilon\bm{\Xi})\mathbf{X}\mathbf{w}\|_{2}^{2}.$
@@
$\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w},\mathbf{y})=-\mathbf{X}^{\top}(\mathbf{C}+\epsilon\bm{\Xi})^{\to}\mathbf{P}\bigl[\mathbf{y}-(\mathbf{C}+\epsilon\bm{\Xi})\mathbf{X}\mathbf{w}\bigr].$
@@
$\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})=-\mathbf{X}^{\top}(\mathbf{C}+\epsilon\bm{\Xi})^{\to}\mathbf{P}\bigl[\mathbf{y}-(\mathbf{C}+\epsilon\bm{\Xi})\mathbf{X}\mathbf{w}^{\star}\bigr].$
@@
$~\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})$
$=$ $~-\underbrace{\mathbf{X}^{\top}(\mathbf{C}+\epsilon\bm{\Xi})^{\top}\mathbf{y}}_{:=\mathbf{T}_{1}}+\underbrace{\mathbf{X}^{\top}(\mathbf{C}+\epsilon\bm{\Xi})^{\top}(\mathbf{C}+\epsilon\bm{\Xi})\mathbf{X}\mathbf{w}^{\star}}_{:=\mathbf{T}_{2}}$
@@
$\frac{\mathrm{d}\mathbf{T}_{1}}{\mathrm{d}\epsilon}=$ $~\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Bigl[-\mathbf{X}^{\top}(\mathbf{C}+\epsilon\bm{\Xi})^{\top}y\Bigr]$
$=$ $~-\mathbf{X}^{\top}\bm{\Xi}^{\top}y.$
@@
$\frac{\mathrm{d}\mathbf{T}_{2}}{\mathrm{d}\epsilon}=$ $~\frac{\mathrm{d}}{\mathrm{d}\epsilon}\Bigl[\mathbf{X}^{\top}(\mathbf{C}+\epsilon\bm{\Xi})^{\top}(\mathbf{C}+\epsilon\bm{\Xi})\mathbf{X}\mathbf{w}^{\star}\Bigr]$
$=$ $~\mathbf{X}^{\top}(\mathbf{C}^{\top}\bm{\Xi}+\bm{\Xi}^{\top}\mathbf{C}+2\epsilon\bm{\Xi}^{\top}\bm{\Xi})\mathbf{X}\mathbf{w}^{\star}$
@@
$\frac{\partial}{\partial\epsilon}\nabla_{\mathbf{w}}\mathcal{L}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X},\mathbf{w}^{\star},\mathbf{y})\Biggr|_{\epsilon=0}=-\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{y}+\mathbf{X}^{\top}(\mathbf{C}^{\top}\bm{\Xi}+\bm{\Xi}^{\top}\mathbf{C})\mathbf{X}\mathbf{w}^{\star}.$
@@
$\mathcal{I}_{\epsilon}(\mathbf{w}^{\star})$ $=-\mathbf{H}^{-1}\Biggl(-\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{y}+\mathbf{X}^{\top}(\mathbf{C}^{\top}\bm{\Xi}+\bm{\Xi}^{\top}\mathbf{C})\mathbf{X}\mathbf{w}^{\star}\Biggr)$
$=\mathbf{H}^{-1}\Bigl(\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr).$
@@
D.3. Edge Influence on Model Output